# Play to the Score: Stage-Guided Dynamic Multi-Sensory Fusion for Robotic Manipulation

### v1

@@ Abstract
Humans possess a remarkable talent for flexibly alternating to different senses when interacting with the environment. Picture a chef skillfully gauging the timing of ingredient additions and controlling the heat according to the colors, sounds, and aromas, seamlessly navigating through every stage of the complex cooking process. This ability is founded upon a thorough comprehension of task stages, as achieving the sub-goal within each stage can necessitate the utilization of different senses. In order to endow robots with similar ability, we incorporate the task stages divided by sub-goals into the imitation learning process to accordingly guide dynamic multi-sensory fusion. We propose MS-Bot, a stage-guided dynamic multi-sensory fusion method with coarse-to-fine stage understanding, which dynamically adjusts the priority of modalities based on the fine-grained state within the predicted current stage. We train a robot system equipped with visual, auditory, and tactile sensors to accomplish challenging robotic manipulation tasks: pouring and peg insertion with keyway. Experimental results indicate that our approach enables more effective and explainable dynamic fusion, aligning more closely with the human fusion process than existing methods.
@@ 1 Introduction
Humans are blessed with the ability to flexibly use various sensors to perceive and interact with the world. Over the years, endowing robots with this potent capability has remained a dream for humanity. Fortunately, with the advancement of computing devices and sensors [1, 2], building a multi-sensory robot system capable of assisting humans has gradually become achievable. We have witnessed the emergence of some exciting multi-sensory robots, such as Optimus and Figure 01.
@@
Figure 1: An illustration for our multi-sensory robot system and a visualization of Modality Temporality in a multi-stage task: pouring. We show confidence in action prediction when using the inputs of all modalities and selectively masking uni-modal features. Due to the changing importance of modalities, both evident inter-stage (coarse-grained) and minor intra-stage (fine-grained) changes in confidence are observed when masking uni-model features. The low confidence fluctuations near stage boundaries also reflect insufficient task stage understanding.
@@ 2 Related Work
Multi-Sensory Robot Learning. Recent works have extensively integrated visual, auditory, and tactile sensors into robot systems. The two common visual modalities, RGB and depth, are widely combined and applied in grasping [16, 17, 18, 19], navigation [20, 21], and pose estimation [22, 23, 24] tasks. In recent years, the development of tactile sensors has sparked interest in the study of tactile modality. Fine-grained tactile data can assist in tasks such as grasping [25, 26, 10, 11], shape reconstruction [27, 28] and peg insertion [29, 13]. Audio modalities are typically applied in tasks such as pouring [9, 12], object tracking [30, 31], and navigation [32, 33, 34]. Some works even attempt to jointly use visual, audio, and tactile modalities to manipulate objects [35, 36, 9]. Based on these efforts, we further explore effective multi-sensory fusion in complex manipulation tasks.
@@
Dynamic Multi-Modal Fusion. Dynamic multi-modal fusion essentially adapts the fusion process based on the input data by assigning different weights to the uni-modal features. A simple and effective way to assign weights is to train a gate network that scores based on the quality of uni-modal features [46, 47, 48, 49]. Another approach to assessing the quality of modalities and dynamically fusing them is uncertainty-based fusion [50, 51, 52]. With the introduction of the transformer [15], self-attention provides a natural mechanism to connect multi-modal signals [53]. Hence, attention-based fusion methods rapidly become a focal point of research [9, 54, 55]. However, these methods primarily perform dynamic fusion based solely on the inputs themselves, while we design dynamic fusion from the perspective of task stages in robotic scenarios.
@@ 3.2 Stage-Guided Dynamic Multi-Sensory Fusion for Robot (MS-Bot)
Stage comprehension module. Merely using the state token alone is insufficient for achieving a comprehensive understanding of the task stages. This module aims to perform coarse-to-grained stage understanding by injecting stage information into the state token. For a task with $S$ stages, we use $S$ learnable stage tokens $[stage_{1}],...,[stage_{S}]$ to represent each stage. Each stage token is initialized with the mean of all state tokens on samples within the stage after warmup. Next, we use a gate network (MLP) to predict the current stage, i.e., coarse-grained stage comprehension. The $S$-dimensional output scores $\mathbf{g}_{t}$ of the gate are softmax-normalized and multiplied by each of the $S$ stage tokens, followed by summing up the results to obtain the stage token $z^{stage}_{t}$ at timestep $t$ via:
$\begin{gathered}\mathbf{g}=(g^{1}_{t},...,g^{S}_{t})=softmax(MLP(z^{state}_{t})),\\ z^{stage}_{t}=\frac{1}{S}\sum_{j=1}^{S}(g^{j}_{t}\cdot[stage_{j}]).\end{gathered}$ (1)
@@
$\begin{gathered}\mathcal{L}_{gate,i}=\sum_{j=1}^{S}(w_{i}^{j}\cdot g_{i}^{j}),\ j\in\{1,2,...,S\},\\ w_{i}^{j}=\left\{\begin{array}[]{ll}1,&(s_{i}=j)\ or\ (\exists k,\ |k-i|\leq\gamma,\ s_{i}\neq s_{k}),\\ 0,&otherwise,\end{array}\right.\\ \end{gathered}$ (3)
@@ 4.1 Task Setup
Fig. 3 briefly shows the two tasks. See Sec. A for a more detailed task setup for both tasks.
@@ 4.2 Physical Setup and Experimental Details
Baselines. We compare our method with three baselines in both tasks: 1) the concat model which directly concatenates all the uni-modal features, 2) Du et al. [59] that uses LSTM to fuse the uni-modal features and the additional proprioceptive information, and 3) MULSA [9] which fuses the uni-modal features via self-attention. We also compare our method with two variants in each task: MS-Bot (w/o A/D) and MS-Bot (w/o T/R) where one modality in the task is removed (audio and touch for pouring while depth and RGB for peg insertion).
Methods Pouring Initial (g) Pouring Target (g) Insertion
@@ 4.3 How much improvement does stage-guided dynamic multi-sensory fusion provide?
Methods Pouring Insertion
@@ 4.5 Can stage-guided dynamic fusion maintain its advantage in out-of-distribution settings?
To verify the generalization of our method to distractors, we conduct experiments with visual distractors in both tasks. In the pouring task, the cylinder’s color was changed from white to red. For the peg insertion task, alterations included changing the base color from black to green (referred to as “Color”), and placing clutter around the base (referred to as “Mess”). As shown in Tab. 3, our method exhibits minimal impact across various scenarios with distractors and consistently maintains performance superiority, demonstrating the generalizability of stage comprehension (Q1, Q3). Two baseline methods, the concat model and Du et al. [59], suffer from performance degradation when the visual modality (including RGB and depth) is disturbed. Due to their lack of ability to understand the stages and dynamically adjust modality weights, even when information from visual modality is not needed in a particular stage, the model is still affected by distractors. This leads to unsatisfying performance. Our method dynamically allocates modality weights based on the understanding of the current stage, thereby reducing the impact of visual distractors on the fused features (Q2, Q3). Consequently, it outperforms the two baseline methods using static fusion and the MULSA that solely relies on the current observation for dynamic fusion.
@@ 5 Conclusion and Limitations
Limitations: The stage division in this work involves human factors. A potential improvement could involve using unsupervised algorithms or leveraging LLMs for stage labeling. Additionally, proposing more challenging multi-sensory manipulation tasks would be interesting for future works.
@@ Appendix C Random Attention Blur
Appendix D Comparison of Attention Scores between MULSA and MS-BOT
@@
Appendix E Detailed Experimental Results of Pouring
Methods Pouring Initial (g) Pouring Target (g)
@@ Appendix E Detailed Experimental Results of Pouring
Methods Pouring Initial (g) Pouring Target (g)
@@
In Sec. 4.4 and 4.5 of the main paper, we show the overall error on all settings of the pouring task. In this section, we comprehensively present the detailed result of the component ablation and the distractor experiment on all the settings of the pouring task in Tab. 4 and 5. As shown in Tab. 4, the contributions from both the state tokenizer and the stage comprehension module on all settings demonstrate the importance of the coarse-to-fine stage comprehension. The consistent lead of MS-BOT across all settings of the pouring task in Tab. 3 also demonstrates the generalizability of stage comprehension.
Appendix F Evaluation of Hyper-parameter Settings

### v2

@@ Abstract
Humans possess a remarkable talent for flexibly alternating to different senses when interacting with the environment. Picture a chef skillfully gauging the timing of ingredient additions and controlling the heat according to the colors, sounds, and aromas, seamlessly navigating through every stage of the complex cooking process. This ability is founded upon a thorough comprehension of task stages, as achieving the sub-goal within each stage can necessitate the utilization of different senses. In order to endow robots with similar ability, we incorporate the task stages divided by sub-goals into the imitation learning process to accordingly guide dynamic multi-sensory fusion. We propose MS-Bot, a stage-guided dynamic multi-sensory fusion method with coarse-to-fine stage understanding, which dynamically adjusts the priority of modalities based on the fine-grained state within the predicted current stage. We train a robot system equipped with visual, auditory, and tactile sensors to accomplish challenging robotic manipulation tasks: pouring and peg insertion with keyway. Experimental results indicate that our approach enables more effective and explainable dynamic fusion, aligning more closely with the human fusion process than existing methods. The code and demos can be found at https://gewu-lab.github.io/MS-Bot/.
@@ 1 Introduction
Humans are blessed with the ability to flexibly use various sensors to perceive and interact with the world. Over the years, endowing robots with this potent capability has remained a dream for humanity. Fortunately, with the advancement of computing devices and sensors [1, 2], building a multi-sensory robot system to assist humans has gradually become achievable. We have witnessed the emergence of some exciting multi-sensory robots, such as Optimus [3] and Figure 01 [4].
@@
Figure 1: An illustration for our multi-sensory robot system and a visualization of Modality Temporality in a multi-stage task: pouring. We show confidence in action prediction (maximum softmax score) when using the inputs of all modalities and selectively masking uni-modal features. Due to the changing importance of modalities, both evident inter-stage (coarse-grained) and minor intra-stage (fine-grained) changes in confidence are observed when masking uni-model features. The low confidence fluctuations near stage boundaries also reflect insufficient task stage understanding.
@@ 2 Related Work
Multi-Sensory Robot Learning. Recent works have extensively integrated visual, auditory, and tactile sensors into robot systems. The two common visual modalities, RGB and depth, are widely combined and applied in grasping [18, 19, 20, 21, 22], navigation [23, 24], and pose estimation [25, 26, 27] tasks. In recent years, the development of tactile sensors has sparked interest in the study of tactile modality. Fine-grained tactile data can assist in tasks such as grasping [28, 29, 12, 13], shape reconstruction [30, 31] and peg insertion [32, 15]. Audio modalities are typically applied in tasks such as pouring [11, 14], object tracking [33, 34], and navigation [35, 36, 37]. Some works even attempt to jointly use visual, audio, and tactile modalities to manipulate objects [38, 39, 11]. Based on these efforts, we further explore effective multi-sensory fusion in complex manipulation tasks.
@@
Dynamic Multi-Modal Fusion. Dynamic multi-modal fusion essentially adapts the fusion process based on the input data by assigning different weights to the uni-modal features. A simple and effective way to assign weights is to train a gate network that scores based on the quality of uni-modal features [49, 50, 51, 52]. Another approach to assessing the quality of modalities and dynamically fusing them is uncertainty-based fusion [53, 54, 55]. With the introduction of the transformer [17], self-attention provides a natural mechanism to connect multi-modal signals [56]. Hence, attention-based fusion methods rapidly become a focal point of research [11, 57, 58, 59]. However, these methods primarily perform dynamic fusion based solely on the inputs themselves, while we design dynamic fusion from the perspective of task stages in robotic scenarios.
@@ 3 Method
In this section, we discuss the challenges in multi-sensory imitation learning from the perspective of task stages and propose the dynamic multi-sensory fusion framework MS-Bot, based on coarse-to-fine task stage understanding, to overcome these challenges. Specifically, the fine-grained state refers to a concept that describes the current situation of the environment, while the coarse-grained stage, divided by sub-goals of the task, is a collection composed of many states, as follows:
$\begin{gathered}\mathbf{E}=\underbrace{(\underbrace{\textbf{X}_{1}}_{State\ 1},a_{1}),(\underbrace{\textbf{X}_{2}}_{State\ 2},a_{2}),...,(\underbrace{\textbf{X}_{t_{1}}}_{State\ t_{1}},a_{t_{1}})}_{Stage\ 1},\underbrace{(\underbrace{\textbf{X}_{t_{1}+1}}_{State\ t_{1}+1},a_{t_{1}+1}),...,(\underbrace{\textbf{X}_{t_{2}}}_{State\ t_{2}},a_{t_{2}})}_{Stage\ 2},...,\end{gathered}$ (1)
where $\mathbf{E}$ is an episode, $\textbf{X}_{t}$ is the multi-sensory observation at timestep $t$ and $a_{t}$ is the action.
@@ 3.2 Stage-Guided Dynamic Multi-Sensory Fusion for Robot (MS-Bot)
Stage Comprehension Module. Merely using the state token alone is insufficient for achieving a comprehensive understanding of task stages. This module aims to perform coarse-to-grained stage understanding by injecting stage information into the state token. For a task with $S$ stages, we use $S$ learnable stage tokens $[stage_{1}],...,[stage_{S}]$ to represent each stage. After warmup training, each stage token is initialized with the mean of all state tokens on samples within the stage. Next, we use a gate network (MLP) to predict the current stage, i.e., coarse-grained stage comprehension. The $S$-dimensional output scores $\mathbf{g}_{t}$ of the gate are softmax-normalized and multiplied by each of the $S$ stage tokens, followed by summing up the results to obtain the stage token $z^{stage}_{t}$ at timestep $t$ via:
$\begin{gathered}\mathbf{g}_{t}=(g^{1}_{t},...,g^{S}_{t})=softmax(MLP(z^{state}_{t})),\\ z^{stage}_{t}=\frac{1}{S}\sum_{j=1}^{S}(g^{j}_{t}\cdot[stage_{j}]).\end{gathered}$ (2)
@@
$\begin{gathered}\mathcal{L}_{gate,i}=\sum_{j=1}^{S}(w_{i}^{j}\cdot g_{i}^{j}),\ j\in\{1,2,...,S\},\\ w_{i}^{j}=\left\{\begin{array}[]{ll}0,&(s_{i}=j)\ or\ (\exists k,\ |k-i|\leq\gamma,\ s_{i}\neq s_{k}),\\ 1,&otherwise,\end{array}\right.\\ \end{gathered}$ (4)
@@ 4.1 Task Setup
Fig. 3 briefly shows the two tasks. See Appendix A for a more detailed task setup for both tasks.
@@ 4.2 Physical Setup and Experimental Details
Baselines. We compare our method with three baselines in both tasks: 1) the concat model which directly concatenates all the uni-modal features [12, 13, 14, 15], 2) Du et al. [63] that uses LSTM to fuse the uni-modal features and the additional proprioceptive information, and 3) MULSA [11] which fuses the uni-modal features via self-attention. We also compare our method with two variants in each task: MS-Bot (w/o A/D) and MS-Bot (w/o T/R) where one modality in the task is removed (audio and touch for pouring while depth and RGB for peg insertion).
Method Pouring-Initial Mass (g) Pouring-Target Mass (g) Insertion
@@ 4.3 How much improvement does stage-guided dynamic multi-sensory fusion provide?
Method Pouring Insertion
@@ 4.5 Can stage-guided dynamic fusion maintain its advantage in out-of-distribution settings?
To verify the generalization of our method to distractors, we conduct experiments with visual distractors in both tasks. In the pouring task, we changed the cylinder’s color from white to red. For the peg insertion task, we altered the base color from black to green (“Color”), and placed clutter around the base (“Mess”). As shown in Tab. 3, our method exhibits minimal impact across various scenarios with distractors and consistently maintains performance superiority, demonstrating the generalizability of stage comprehension (Q1, Q3). Two baseline methods, the concat model and Du et al. [63], suffer from performance degradation when the visual modality is disturbed. Due to their lack of ability to understand the stages and dynamically adjust modality weights, even when information from visual modality is not needed in a particular stage, the model is still affected by distractors. This leads to unsatisfying performance. Our method dynamically allocates modality weights based on the understanding of the current stage, thereby reducing the impact of visual distractors on the fused features (Q2, Q3). Consequently, it outperforms the two baseline methods using static fusion and the MULSA that solely relies on the current observation for dynamic fusion.
@@ 5 Conclusion and Limitations
Limitations: Our stage division involves human factors. An improvement could be leveraging LLMs for automatic stage labeling. Additionally, proposing more challenging multi-sensory manipulation tasks would be interesting for future works. More limitations are discussed in Appendix L.
Acknowledgments
The project was supported by National Natural Science Foundation of China (NO.62106272).
@@ Appendix C Random Attention Blur
Appendix D LLM-Based Automatic Stage Labeling
Our strategy for stage labeling is to segment the action sequence of the trajectory based on pre-set rules (such as considering the timestep of the first gripper rotation as a stage transition point). This does indeed require some time to pre-define the stages and design specific rules for division. However, extracting stage information without human annotation is also feasible. One possible approach is to use LLMs, which have demonstrated strong comprehension ability in robotic manipulation [66], for automatic stage annotation of action sequences. We use Claude 3.5 Sonnet [67] as our LLM, and construct prompts to describe the task setup and include an example action sequence to better help the LLM understand the process of this task. We then ask the LLM to provide the specific stage division, as shown in Fig. 9.
Figure 9: An example of automatic stage division by LLM. We provide the task setup, robot action space and an example action sequence to the LLM to obtain the main stage division. For ease of presentation, the action sequence in the figure has been abbreviated.
After obtaining the stage divisions provided by the LLM, we then construct another prompt and ask the LLM to label the action sequence according to the stage divisions, as shown in Fig. 10. This strategy allows for obtaining stage labels for all samples without requiring human annotation. We find that the stage divisions provided by the LLM are usually more detailed. For example, it divides the peg insertion with keyway task into four stages (splitting our First Insertion stage into two stages: Initial Positioning and Vertical Approach), which is also reasonable. Since our model does not explicitly require that the modality importance must be different between stages, a more detailed stage division is acceptable to some extent.
Figure 10: An example of automatic stage labeling by LLM. We provide the task setup, robot action space and an target action sequence to the LLM to obtain the stage labels. For ease of presentation, the action sequence in the figure has been abbreviated.
Through this method, we perform automatic stage annotation on our imitation learning dataset. To evaluate the accuracy of the LLM’s annotations, we merge the first two stages divided by the LLM into a single First Insertion stage, and use the human-annotated stage labels as ground truth. The evaluation results show that the accuracy of the stage labels annotated by the LLM is $95.68\%$ (10,096 / 10,552, about 15 errors per trajectory).
For tasks like peg insertion with keyway, where many actions are continuous, we can merge consecutive actions to obtain a compressed action sequence like “[action*num,…,action*num]”, as shown in Fig. 11. This approach helps reduce annotation errors caused by LLM counting mistakes and increases the accuracy to $99.65\%$ (10,521 / 10,552, about 3 errors per trajectory). These results demonstrate that automatic stage annotation using LLMs without pre-defined stages is feasible. The experiment is only a relatively simple attempt. We believe that with more detailed prompt engineering and the use of more powerful LLMs, the accuracy can be further improved.
Figure 11: An example of automatic stage labeling by LLM. We provide the task setup, robot action space and an compressed target action sequence to the LLM to obtain the stage labels.
Appendix E Comparison of Attention Scores between MULSA and MS-Bot
@@
Appendix F Combination with Diffusion Policy
Fusion Module Policy Network Success Rate
Concat MLP 7/15
Concat Diffusion Policy [68] 11/15
Stage-guided Dynamic Fusion MLP 13/15
Stage-guided Dynamic Fusion Diffusion Policy [68] 13/15
Table 4: Comparison of performance with Concat and Diffusion Policy baselines on the peg insertion with keyway task.
To verify the applicability of our proposed stage-guided dynamic multi-sensory fusion to different policy networks, we replace the policy network with the Diffusion Policy [68] and conduct experiments on the peg insertion with keyway task. Specifically, we replace the MLP policy network with an UNet-based Diffusion Policy in the Concat baseline and our MS-Bot model, and evaluate the performance of stage-guided dynamic fusion on models using different policy networks.
The results in Tab. 4 show that replacing the policy network in the Concat baseline with the Diffusion Policy leads to a significant performance improvement, but there is still a gap compared to MS-Bot. This is because the Diffusion Policy still lacks an explicit feature fusion module (apart from simple concatenation). The fusion of multi-sensory features in this baseline still occurs within the policy network, similar to the concat baseline, albeit with a more capable policy network. After further replacing its fusion module with our proposed stage-guided dynamic fusion, it achieves performance comparable to MS-Bot. This demonstrates that our proposed explicit stage understanding and stage-guided dynamic fusion are still effective for models with more powerful policy networks. Our proposed fusion module can help the policy network focus more on generating actions using the already fused multi-sensor features. However, under the same conditions of using stage-guided dynamic fusion, the model with diffusion policy as the policy network does not outperform the model with MLP. We think this may be due to the relatively small scale of data in this task, which does not allow the Diffusion Policy to fully leverage its advantages.
Appendix G Comparison with Stage-Aware and Context-Aware Baselines
Method Success Rate
Concat 7/15
MULSA 11/15
Concat + Stage Auxiliary Task 9/15
MULSA + Stage Auxiliary Task 11/15
MULSA + Observation Context 11/15
MS-Bot 13/15
Table 5: Comparison of performance with stage-aware and context-aware baselines on the peg insertion with keyway task. We introduced stage prediction auxiliary tasks and multi-sensory observation context into the Concat and MULSA baselines for a more comprehensive comparison.
Due to not fully considering the importance of explicit stage understanding in guiding multi-sensory fusion, the original Concat and MULSA baselines does not utilize contextual inputs and additional stage information. To make a fairer comparison with our MS-Bot, we respectively construct Concat and MULSA baselines that incorporate a stage prediction auxiliary task and multi-sensory observation context. Specifically, we add an extra head (an MLP identical to the gate network in MS-Bot) to both the Concat and MULSA models to predict the stage, trained with the same form of loss used to penalize the scores of the gate network in MS-Bot. We also attempt to modify the structure of the MULSA model to accept observation context as input. We input the observations from 5s, 7.5s, 10s, 12.5s, 15s, 17.5s, and 20s before the current timestep into the model, and encode these observation contexts by each modality’s encoder into 7 tokens (per modality). These tokens are then concatenated with the original 6 tokens of the current timestep observation of each modality, resulting in a total of 13 tokens for each modality. We evaluate these baselines on the peg insertion with keyway task. The results in Tab. 5 shows that introducing an explicit stage prediction task could improve the performance of the Concat baseline, but it provides minimal benefit for the MULSA method. These baselines still have a performance gap compared to our MS-Bot. This indicates that explicit stage understanding is meaningful, but introducing context to aid stage understanding and specially designed stage-guided dynamic fusion modules are also necessary. In addition, the results show that inputting observation contexts into MULSA results in virtually no improvement in the model’s performance. This indicates that mixing the contexts and current observations in self-attention calculations is highly inefficient for helping the model understand stages and allocate modality weights, as it reduces focus on the more important current observations and significantly increases GPU memory usage. In contrast, our MS-Bot leverages a long action history as context to enhance its understanding of task stages. This approach simplifies processing for the model and offers finer granularity, effectively guiding dynamic multi-sensor fusion.
Appendix H Comparison under the Multi-Camera Setting
Method Success Rate
Concat 9/15
MULSA 12/15
MS-Bot 14/15
Table 6: Comparison under the setting of multiple cameras on the peg insertion with keyway task.
In robotic manipulation, a common setup involves using multiple cameras to capture visual information from different views. By treating visual images from different views as distinct modalities, our MS-Bot is also able to handle multi-camera observations. To verify this capability, we conduct a comparison with baseline methods on the peg insert with keyway task under the setting with multiple cameras. Specifically, we place another camera on the side of the base, so that the lines connecting these two cameras to the base roughly form a right angle. We replace the depth observation from the first camera with the RGB observation from the second camera, as the RGB observations from these two cameras already provide sufficient spatial information for this task. The results in Tab. 6 show that in the multi-camera setup, our MS-Bot method still outperforms the baselines. This demonstrates that our proposed stage-guided dynamic fusion can also be applied to fuse features from multiple cameras.
Appendix I Combination with Audio Modality
Method Success Rate
Concat 8/15
MULSA 11/15
MS-Bot 13/15
Table 7: Comparison on the peg insertion with keyway task using RGB, depth, tactile and audio.
For tasks involving using lock-picking tools to unlock in real life, audio can indeed serve as the basis for completing the task, as these locks typically have specific mechanical structures that produce distinct sounds. Similarly, audio can indicate whether the key and keyway are aligned, thereby assisting with the peg insertion with keyway task. To verify the effectiveness of our model after further incorporating audio, we conduct experiments on the peg insertion with keyway task using RGB, depth, tactile and audio modality. The results in Tab. 7 show that the addition of the audio modality only brings improvement to the weakest Concat baseline. For MULSA and our MS-Bot method, the benefits are minimal. This is because the audio in this task is short and minimal due to the 3D-printed materials and is vulnerable to noise disturbance.
Appendix J Detailed Experimental Results of Pouring
Method Pouring Initial (g) Pouring Target (g)
@@ Appendix I Combination with Audio Modality
Method Pouring Initial (g) Pouring Target (g)
@@
In Sec. 4.4 and 4.5 of the main paper, we show the overall error on all settings of the pouring task. In this section, we comprehensively present the detailed result of the component ablation and the distractor experiment on all the settings of the pouring task in Tab. 8 and 9. As shown in Tab. 8, the contributions from both the state tokenizer and the stage comprehension module on all settings demonstrate the importance of the coarse-to-fine stage comprehension. The consistent lead of MS-BOT across all settings of the pouring task in Tab. 3 also demonstrates the generalizability of stage comprehension.
Appendix K Evaluation of Hyper-parameter Settings
@@
Appendix L Limitations and Future Improvements
In this section, we discuss the potential limitations of our work and the corresponding solutions:
In many robotic tasks this kind of dynamic weighting is not necessary. Indeed, there are many tasks that can be completed with just the visual modality, where readings from other sensors may not provide useful information. We believe that one solution is to let the model first select the necessary subset of sensors based on the task (possibly using an LLM), and then perform dynamic multi-sensor fusion based on stage understanding. For tasks with only one stage, this can be considered as a special case where $|S|=1$.
The stage labeling may be costly. Currently, our strategy for stage labeling is to segment the action sequence of the trajectory based on pre-set rules. This requires time to set up detailed rules and becomes more challenging when there are more stages and more complex stage transitions. A possible solution is the LLM-based automatic labeling method mentioned earlier, although a small amount of human effort is still necessary, such as constructing prompts to accurately describe the task setup and providing example for in-context learning.
The lack of comparison to larger models. Currently, there is no well-trained open-source multi-sensory large foundation model for robot manipulation (including vision, hearing, touch, etc.). Therefore, we are temporarily unable to verify the role of explicit stage understanding in large models, nor can we directly compare our model with multi-sensory large models. However, we believe that even for larger models, explicit stage understanding remains meaningful when introducing new modalities and training is required using a small amount of paired multi-sensory data, or when fine-tuning in new scenarios with limited data scale.
Robustness issue in the case of missing modalities. Since in both of our tasks there are stages that depend on a single modality, masking uni-modal features during testing may directly lead to failure. This is a problem common to all multi-sensory models, and we did not make improvements specifically for this issue. In cases where multiple sensors complement each other, the absence of uni-modal features may have a compounding effect on stage understanding and dynamic fusion. We believe that one possible solution is to randomly dropout uni-modal features during the training process [69]. Some other methods for robust multi-modal learning could also be referenced.
Transferability of the current dynamic multi-sensory fusion architecture. Since the multi-sensory fusion architecture we use in the dynamic fusion module is based on cross-attention rather than self-attention, there are difficulties in directly transferring the original form of this fusion method to transformers with multiple self-attention blocks. One potential strategy is to use the stage token as a special token to replace the class token. It could be used to predict the stage under the supervision of stage labels and thereby achieving the effect of stage-guided multi-modal fusion.
Low integration level of current multi-sensory robot system. Our current multi-sensory robot system has not well integrated the various sensors: the camera is positioned at a third-person perspective, the microphone is placed on the desktop, and the tactile sensor is at the end of the robotic arm. They are located in different positions. This low level of integration means that the system can only complete relatively simple tasks, while more complex multi-sensory tasks that require mobility cannot be accomplished. This multi-sensory system still needs further refinement and upgrading.